OpenStudy (yuii):

What is the common ratio for the geometric sequence? 54, 36, 24, 16...

OpenStudy (marcelie):

formula = 1 ___________ 1-r i believe that you would get your numbers by subtracting them and that would be your r

OpenStudy (chrishogan):

So @marcelie You're saying r=-22??

OpenStudy (mathstudent55):

The common ratio is the number you multiply a term to find the next term. To find the common ratio, pick any two consecutive numbers, and divide the one term by the previous term.

OpenStudy (mathstudent55):

In your case, you can find the common ratio by doing any of these divisions: \(\dfrac{36}{54} \) \(\dfrac{24}{36} \) \(\dfrac{16}{24} \) All of them will give you the same answer.

OpenStudy (yuii):

Im confused..

OpenStudy (mathstudent55):

What are you confused with? In a geometric sequence, if you divide any term by the previous term, you find the common ratio. Since you have 4 given terms, you can find the common ratio using one of the 3 fractions I gave you above. You don't need to do all three fractions. They all reduce to the same fraction. Just reduce one of the fractions to lowest terms.